### v1

# Towards Comprehensive Multimodal Perception: Introducing the Touch-Language-Vision Dataset

Abstract
Tactility provides crucial support and enhancement for the perception and interaction capabilities of both humans and robots. Nevertheless, the multimodal research related to touch primarily focuses on visual and tactile modalities, with limited exploration in the domain of language. Beyond vocabulary, sentence-level descriptions contain richer semantics. Based on this, we construct a touch-language-vision dataset named TLV (Touch-Language-Vision) by human-machine cascade collaboration, featuring sentence-level descriptions for multimode alignment. The new dataset is used to fine-tune our proposed lightweight training framework, TLV-Link (Linking Touch, Language, and Vision through Alignment), achieving effective semantic alignment with minimal parameter adjustments (1%). Project Page: https://xiaoen0.github.io/touch.page/.
1 Introduction
Tactile perception occupies a distinctive and pivotal role within the human sensory system, constituting a fundamental basis for our cognitive comprehension of the environment, coexisting harmoniously with other sensory modalities, such as vision and audition. Tactility allows us to perceive the texture, temperature, and hardness of objects etc., and enables us to explore environments and perform intricate tasks, such as grasping and manipulating. The significance of touch is evident not only in humans Dahiya et al. (2009); Johansson and Flanagan (2009) but also in robotic applications Qi et al. (2023); Hansen et al. (2022), where the acquisition and processing of tactile information are crucial for enhancing the perceptual capabilities and interaction efficiency of these applications.
Despite the undeniable significance of touch, tactile-related multimodel research predominantly focuses on the visual and tactile modalities Dave et al. (2024); Yang et al. (2023); Kerr et al. (2022), with limited exploration in the domain of language. While there are some works related to language, they remain primarily at the lexical level, serving as labels for classification purposes Yang et al. (2022); Gao et al. (2020); Yuan et al. (2017a). This arises from the heightened challenges associated with annotating lengthier texts, including intricate narratives and elevated expenses.
Figure 1: Construction process of the TLV dataset.
Continuous innovation in image-to-text models Achiam et al. (2023); Team et al. (2023) enables the generation of fluent text from prompts and images, thereby offering opportunities for tactile annotation with longer texts. In this work, we introduce a tactile-related multimodel dataset, named TLV (Touch-Language-Vision), through human-machine cascade collaborative annotation. TLV incorporates three modalities: touch, language, and vision, with pairwise correspondence between any two modalities, aiming to strengthen alignment between touch and language. Compared to a set of vocabularies (i.e., lexical-level descriptions), the descriptions in TLV are at the sentence level, capable of conveying more rich and more complete semantic information.
To assess TLV’s efficacy, we employ it as the training dataset and present a lightweight unsupervised training method, TLV-Link (Linking Touch, Language, and Vision through Alignment). This method maps all modalities to a shared embedding space, enabling effective semantic alignment. To improve training efficiency, we employ Low-Rank Adaptation (LoRA) Hu et al. (2021) for fine-tuning and only 1% of the parameters are adjusted. Subsequently, we evaluate the performance of TLV-Link on various tactile classification tasks using a cross-domain dataset. Experimental results demonstrate the potential of the TLV dataset. This paper presents the following contributions:
  • Introducing TLV, a new touch-language-vision dataset with sentence-level descriptions annotated by human-machine cascade collaboration, Addressing the challenge of tactile annotation for longer texts.
  • Proposing TLV-Link, a lightweight joint pretraining framework characterized by independence from labeled data, the utilization of a smaller dataset, the adjustment of model parameters, and acceptable performance.
  • Validating the effectiveness of our dataset and method and providing direction for further optimization on tactile-related tasks.
2 Related Work
Tactile Perception
Extracting and leveraging tactile information, encompassing surface texture, elasticity, and temperature, holds substantial promise for advancements in both robotics and AI research Cui et al. (2020); Fazeli et al. (2019); Lin et al. (2019). Current tactile sensors primarily rely on vision, employing a camera and illumination system to record deformations in a curved elastomeric gel. This structure has given rise to diverse perception systems, including GelSight Yang et al. (2022); Li et al. (2019); Johnson et al. (2011); Gomes et al. (2021); Calandra et al. (2017); Yuan et al. (2017b); Si and Yuan (2022); Dong et al. (2017), DIGIT Kerr et al. (2022); Suresh et al. (2023); Lambeta et al. (2020), Taxim Gao et al. (2022), GelSlim Gao et al. (2023) and TACTO Gao et al. (2021). These systems aim to comprehensively record high-resolution, detailed tactile information. Among them, GelSight Yang et al. (2022) stands out as one of the most widely used tactile perception systems, offering elaborate capture of depth, shear, and surface orientation. This work leverages sensor images primarily from GelSight.
Tactile Datasets
A significant challenge in learning from the tactile modality lies in the substantial human effort and time required to construct high-quality datasets. Despite this hurdle, the research community’s continuous efforts have yielded several publicly available datasets: Objectfolder 2.0 Gao et al. (2021) (featuring 1,000 implicitly represented objects generated through simulation), SSVTP Kerr et al. (2022) (containing 4.5K spatially aligned image-tactile pairs acquired using DIGIT), the Feeling of Success Calandra et al. (2017) (employing a two-finger gripper and GelSight sensor), Touch and Go Yang et al. (2022) (a high-quality, in-the-wild dataset encompassing diverse categories and quantitative visuo-tactile pairs) and VisGel Li et al. (2019) (a dataset comprising over 12K touch instances and 3 million vision-touch frames). However, these datasets are in the absence of rich textual descriptions, hindering their potential for realizing higher-level cross-modal alignment. This work addresses this limitation by incorporating detailed and qualitative captions, fostering a more comprehensive and advanced cross-modal understanding.
Multimodal Alignment
Effectively aligning semantics from diverse modalities is fundamental and crucial in multimodal research, yet it had been challenging to construct a high-dimensional joint embedding space incorporating features of different modalities. CLIP Radford et al. (2021) achieved remarkable performance and generalization ability through self-supervised contrastive pretraining on a massive dataset of 400 million image-text pairs scraped from the internet. Subsequent works such as ALIGN Jia et al. (2021), Flamingo Alayrac et al. (2022), Open-CLIP Schuhmann et al. (2022) have further advanced the field towards more robust and accurate alignment. Beyond vision and language modalities, significant research efforts aim to bridge the gap between even more diverse modalities, including 3D points Liu et al. (2024); Xue et al. (2023); Zhang et al. (2022); Guo et al. (2023), audio Guzhov et al. (2022); Chen et al. (2023). ImageBind Girdhar et al. (2023) significantly extended the joint embedding space to encompass six distinct modalities through image-centered contrastive learning, further promoting comprehensive cross-modal understanding. Along this line, LanguageBind Zhu et al. (2023) proposed a language-centered alignment strategy to fully leverage the rich semantic information within the text, achieving significant performance improvements. This work builds upon these advancements by further expanding cross-modal alignment to concurrently include touch alongside other modalities.
3 TLV Dataset
The TLV dataset aims to associate tactile and visual perceptions with sentence-level descriptions for multimodal alignment. As shown in Figure 1, the construction process of TLV consists of three stages: touch and vision collection (Sec. 3.1), touch localization (Section. 3.2), and tactile labeling (Sec. 3.3).
3.1 Stage I: Touch and Vision Collection
We collect paired tactile and visual observations from VisGel Li et al. (2019), a large vision-touch dataset collected by a video camera, and a tactile sensor called GelSight Johnson et al. (2011). VisGel captured synchronized videos of the scenes where the robotic arm touched the objects and recorded timestamps to synchronize visual and tactile images. Among the synchronized videos captured, 10,000 videos were used to construct the training dataset. We utilize the synchronized visual and tactile images from these 10,000 synchronized videos for touch and vision collection.
From the visual videos, we observe that the first frame depicts the starting state of the robotic arm when it is away from objects. As time progresses, the arm gradually approaches an object until it makes contact, remains in contact for a period, and then slowly withdraws. Based on the above observations, for each pair of synchronized videos, we select two sets of synchronized visual and tactile frames: one set depicting an object being touched, and the other set showing no object being touched. To obtain frames where an object is being touched, we use the first frame as the background and apply frame differencing Zaki et al. (2011). The frame with the maximum difference from the background is selected as the frame where the object is being touched. Through observation, we uniformly select the 40th frame as the frame where no object is being touched.
3.2 Stage II: Touch Localization
The two modalities of touch and vision can be regarded as different views containing the same semantics. From this standpoint, we recruit participants to label the object being touched in visual images from Stage I. For visual images where no object is touched, we do not consider them. This serves as preliminary touch localization, preparing for the next step of tactile labeling using GPT-4V Achiam et al. (2023). touch localization comprises two parts: highlighting the touched object with a red box in visual images and providing a name for the enclosed object. The labeling of object name is open-ended, and we do not provide a predefined set of candidate object names. In the process of object labeling, we found that due to issues during the collection of the original dataset (i.e. VisGel), certain data could not be annotated. For example, instances where the touched object is occluded or the entire video does not involve interaction with any object. We have filtered out such data.
3.3 Stage III: Tactile Labeling
From the perspective of containing identical semantics in both touch and vision, we utilize GPT-4V for the annotation of texts. For each visual image with the highlighted box from Stage II, we employ a thoughtfully designed, data-specific prompt. This prompt instructs GPT-4V to generate detailed descriptions, taking into account factors such as the name of the touched object, the specific location of the contact, the material composition at the point of contact, and the texture characteristics and softness/hardness of the touched area. For visual images where no object is touched, we refrain from using GPT-4V for annotation and instead provide an uniform description: No object is being touched..
3.4 Dataset Statistics
We have annotated text-based descriptions for 20,000 pairs of synchronized tactile and visual observations collected from VisGel, including 10,000 pairs with an object being touched and 10,000 pairs without an object being touched. For the cases where an object is touched, we filtered out data that cannot be annotated as mentioned in Stage II, resulting in the annotation of 9,843 instances. For the cases without an object being touched, we annotated all 10,000 instances. Thus, we ultimately obtained a total of 19,834 annotated data entries. To our knowledge, this is the first touch-language-vision dataset with sentence-level descriptions.
Figure 2: Overview of our lightweight joint training method.
4 Method
We propose TLV-Link (Linking Touch, Language, and Vision through Alignment), an unsupervised and lightweight joint training method designed to leverage the TLV dataset we constructed. The visual observations in TLV can be considered as auxiliary information, assisting in learning the alignment between touch and language and enhancing the zero-shot classification ability of touch. The method primarily consists of three components: multi-modal encoder, LoRA fine-tuning, and joint training, as illustrated in Fig. 2.
4.1 Multi-modal Encoders
TLV-Link involves three modalities: touch, language, and vision. We treat the touch modality as RGB images for processing. Therefore, for both touch and vision modalities, we use the Vision Transformer (i.e., ViT) Dosovitskiy et al. (2020) for encoding. The touch and vision encoders are instantiated as OpenCLIP Ilharco et al. (2021) vision encoders. For the text encoder, we instantiate it as an OpenCLIP text encoder.
4.2 LoRA Fine-tuning
Differing from prior approaches Lei et al. (2023), we do not utilize large-scale datasets for pre-training. Instead, we employ LoRA Hu et al. (2021) for lightweight fine-tuning on the TLV dataset. For a modality-agnostic encoder $f(\cdot)$ with a weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, we maintain the weight matrix $W_{0}$ frozen while learning a new weight matrix $BA$. The forward pass can be formalized as follows:
$f({\displaystyle{\bm{x}}})=W_{0}{\displaystyle{\bm{x}}}+BA\bm{x}$ (1)
where $B\in\mathbb{R}^{d\times r},A\in\mathbb{R}^{r\times k}$, with $r$ being the minimum of $d$ and $k$.
4.3 Joint Training
Joint learning aims to align touch and language better. While learning the alignment between touch and language, we also acquire knowledge about the alignment between vision and language, as well as the alignment between touch and vision. The text encoder from OpenCLIP has demonstrated good generalization in text, so during the joint learning process, we freeze the text encoder and only update the touch encoder and vision encoder. The update to the vision encoder is made to assist the update to the touch encoder. To ensure alignment across different modalities, we perform contrastive learning principles Radford et al. (2021) for joint learning.
$L_{T,L}=-\frac{1}{K}\sum_{i=1}^{K}\log\frac{\exp(x_{i}^{\top}y_{i}/\tau)}{\sum_{j=1}^{K}\exp(x_{i}^{\top}y_{j}/\tau)},$ (2)
$L_{V,L}=-\frac{1}{K}\sum_{i=1}^{K}\log\frac{\exp(z_{i}^{\top}y_{i}/\tau)}{\sum_{j=1}^{K}\exp(z_{i}^{\top}y_{j}/\tau)},$
$L_{T,V}=-\frac{1}{K}\sum_{i=1}^{K}\log\frac{\exp(x_{i}^{\top}z_{i}/\tau)}{\sum_{j=1}^{K}\exp(x_{i}^{\top}z_{j}/\tau)}$
where $x,y,z$ represent the observations of tactile, language, and visual modalities, respectively, and $\tau$ and $K$ are the scalar temperature and batch size. In practice, we use a symmetric joint loss $(L_{T,L}+L_{L,T})+\alpha(L_{V,L}+L_{L,V})+\beta(L_{T,V}+L_{V,T})$.
Model Size Touch and Go
Material Hard/Soft Rough/Smooth
ImageBind Base 24.2 65.7 69.8
V-L-2 (I) Base 29.9 (+5.7%) 72.4 (+6.7%) 77.9 (+8.1%)
V-L-2 (I) Large 31.2 (+7.0%) 74.3 (+8.6%) 78.2 (+8.4%)
V-L-2 (I+T) Large 65.8 (+41.6%) 74.7 (+9.0%) 63.8 (-6.0%)
OpenCLIP Large 17.7 32.2 42.7
TLV-Link Large 26.0 (+8.3%) 65.1 (+32.9%) 74.6 (+31.9%)
Table 1: Accuracy of different models on various tactile classifications. Those with performance improvement compared with their respective foundation within 30% are marked as green and above 30% are marked as red. V-L-2: VIT-LENS-2; (I): Anchored by images; (I+T): Anchored by images and texts.
5 Experiments
5.1 Setup
We evaluate our model and dataset on various tactile classification tasks, including material, hard/soft, and rough/smooth classification, using the Touch-and-Go dataset Yang et al. (2022). This means that zero-shot evaluation is conducted on a cross-domain dataset. TLV-Link is extended based on OpenCLIP-large Ilharco et al. (2021) and fine-tuned on our TLV dataset in an unsupervised and lightweight manner. Because the visual modality is considered as auxiliary information, both $\alpha$ and $\beta$ are set to 0.1 in the symmetric joint loss. We use accuracy as the metric.
5.2 Results and Analysis
We contrast our model with ViT-LENS-2 Lei et al. (2023), a state-of-the-art multi-model that excels in zero-shot performance on tactile tasks. The comparative results of different models can be found in Table 1. While the accuracy of TLV-Link may be not optimal, exhibits an 8.3% improvement in material classification compared to our foundation, OpenCLIP. Especially significant is the marked improvement in both hard/soft and rough/smooth classifications, with TLV-Link’s performance advancing by more than 30%. Nevertheless, VIT-LENS-2 (I) demonstrated an improvement ranging from 7% to 9% compared to their foundation, ImageBind Girdhar et al. (2023). Despite VIT-LENS-2 (I+T) displaying a notable 41.6% boost in material classification, there was a 6% decline in rough/smooth classification. This reflects the effectiveness of the TLV dataset and the efficiency of TLV-Link in utilizing data. Certainly, we recognize certain performance limitations of TLV-Link, thus prompting us to analyze distinctions in training paradigm, #training data, parameter tuning ratio, and cross-domain evaluation between TLV-Link and ViT-LENS-2, as illustrated in Table 2. It can be observed that TLV-Link is characterized by independence from labeled data, the utilization of a smaller dataset, a lightweight training approach, and evaluation across various data domains. This could make it more attractive for specific application scenarios.
Model TP #TD PTR CDE
VIT-LENS-2 Supervised 91,982 100%
TLV-Link Unsupervised 19,843 1%
Table 2: Comparison of Ours and VIT-LENS-2 in training paradigm (TP), #training data (#TD), parameter tuning ratio (PTR), cross-domain evaluation (CDE).
5.3 Ablation Study
We conduct the ablation study in Table 3 to illustrate the impact of vision information. Simultaneously aligning both touch and text with visual information enhances tactile classification, yielding a positive overall effect. Conversely, aligning either touch or text with visual information has a detrimental effect.
Model Touch and Go
Material Hard/Soft Rough/Smooth
TLV-Link 26.0 65.1 74.6
-TV 27.8 52.8 52.7
-VL 26.5 55.3 49.1
-(TV&VL) 32.5 56.5 56.6
Table 3: Impact of vision information at different levels. -TV: Do not align touch with vision; -VL: Do not align language with vision; -(TV&VL): Do not involve visual information.
6 Conclusion
In this work, we construct the first touch-language-vision dataset, TLV, featuring sentence-level descriptions for multimodal alignment. To demonstrate the effectiveness of the TLV dataset, we extended OpenCLIP and proposed TLV-Link, an unsupervised lightweight training approach. Preliminary experiments validate that the TLV dataset facilitates better alignment between touch and language. The proposed method may apply to specific scenarios, but there is room for improvement in terms of performance, and further enhancements are needed. Additionally, we intend to extend the application of TLV to more tasks to fully exploit its potential.